# Efficient system for bulk characterization of cryogenic CMOS components

### v1

@@ I Introduction
Quantum computing devices based in semiconductor systems offer many possible advantages of ready scalability by harnessing existing fabrication technology [1]. Application-specific integrated circuits (ASICs) fabricated using established foundry processes can be operated at low temperature to support the scalable control of the large number of qubits necessary for error correction [2]. Integrated multiplexing and signal processing reduces the number of connecting lines between cryogenic stages and room-temperature instruments and reduces the complexity of the room-temperature readout instrumentation required [3].
At the most fundamental level, confident cryogenic integrated circuit design begins with test and validation of the elementary transistor components in the technology node under realistic cryogenic operation conditions. This is necessary because the operating parameters of silicon devices are dependent on temperature, and this dependence cannot be extrapolated from foundry-provided models (often only rated down to -55 °C [4]) due to the onset of cryogenic effects which contribute significantly to very low-temperature behavior. The device parameters captured in the process design kit (PDK) used in electronic design automation (EDA) tools need to accurately reflect realistic operation at low temperature.
@@
Capturing a comprehensive set of cryogenic test data presents logistical challenges. The overhead of wire-bonding individual transistor device terminals combined with practical constraints of die size, cryostat wiring, and thermal cycle times make this approach impractical. At room temperature, these measurements are typically performed at the wafer scale, using an automated wafer prober to access dedicated pads for each transistor across large arrays of devices. There are advantages to this approach. Wafer-level probing of individual devices minimizes the de-embedding and leakage compensation efforts required due to minimal device layout parasitic elements and the absence of any active devices in the measurement path. Particularly in the case of dynamic behavior characterization, this approach is invaluable to ensure measurement accuracy. It has been shown that this method can be extended to cryogenic environments by utilizing highly specialized cryogenic probe systems [6]. This method allows wafer-scale variability to be measured in an analogous way to the traditional foundry PDK development process; however, it lacks flexibility and accessibility due to the heavy dependence on expensive, specialist equipment and inefficient use of wafer area.
@@ III Measurement System
Figure 1: (a) Block diagram of the key components in the measurement system. A die containing multiple device arrays, each consisting of 1024 transistors accessed via a multiplexer is wire-bonded to a demount-able PCB "daughter card." A corresponding multi-pin socket mounted to a He dipping probe provides an interface to the digital circuitry necessary to configure the state of the multiplexer as well as the analog circuity used to conduct the measurements on the selected device. (b) Photograph of the bonded die. By recessing the die within a multi-level PCB and utilizing the surface of both levels, the density of bond pads can be increased, minimizing the length and approach angle of bond wires needed.
@@ III-B Integrated Circuits (DUTs) and Sample Packaging
The core multiplexer unit cell structure contains an analog signal pass gate formed by an N-channel FET (NFET) and P-channel FET (PFET) in parallel with digitally driven gate voltages to define the on and off states as shown in Figure 2(a). Since the increased threshold voltages of these devices at cryogenic temperatures result in substantially increased on-resistance when the pass voltage is near the midpoint of the high and low supply voltages, $V_{DD}$ and $V_{SS}$ (in some cases both the NFET and PFET may be in the subthreshold regime even with the pass gate in the "on" state), the backgate voltages (not shown) of the multiplexer pass gate FETs are externally driven to enable shifting the device thresholds nearer to their room-temperature values [11].
@@ III-E Analog Instrumentation
Figure 2: (a) High-level circuit diagram of an on-die multiplexer structure allowing analogue connections to be established with Kelvin sense to each terminal of a selected transistor. Gate contacts are established at opposite sides of the channel to allow gate resistance thermometry to be performed. A pair of 32-bit registers (one for one-hot row selection and another for one-hot column selection) are used to turn on the pass gate to the device at the selected row/column pair. Each signal pass gate is formed by an NFET and PFET in parallel. (b) Diagram of analog instrument connections used to conduct a 6-terminal measurement with Kelvin sense (FH, SH, FL, and SL corresponding to force-high, sense-high, force-low, and sense-low channels of SMUs, respectively). The ESD protection diodes in blue demonstrate the value of using SMU3 as a reference to the system ground. SMU3 in this case can bias the source at an arbitrary reference potential with respect to system ground, enabling full control and optimization of measurements to avoid excess forward-bias currents flowing in any ESD diodes while maintaining high-impedance sense lines.
Analog transistor DC characteristics can be probed with multiple source-measure units (SMUs) as shown in Figure 2(b). Typically, one unit controls drain-source bias and the other controls gate-source bias. These are bound to a common reference point whose potential can be arbitrarily set with respect to other chassis ground by the third SMU. When using a Kelvin sensing configuration, each device terminal must have independent lines for current injection ("force") and voltage monitoring ("sense"). These can be implemented as parallel switching routes to a multiplexed device terminal as previously described. Referring voltage sources to locally sensed voltages on the device side of the multiplexers is essential due to additional series resistances in measurement leads and in multiplexer signal paths. This system also allows the analog circuitry as a whole to be referenced to the ESD protection low-side supply $V_{SS}$ in a dynamic way, allowing the full range of accessible potentials defined by ESD protection structures to be utilized regardless of the polarity of relative potentials to be applied to the device under test (DUT).
@@ III-F Leakage Subtraction
Leakage on the die arises mainly from ESD protection structures, off-state drain-source leakage through the disabled pass FETs of other multiplexer channels, off-state leakage through the disabled pull-down/-up FETs, and gate leakage from the various FETs. These leakage effects can be measured (and then subtracted from actual device measurements) by measuring the open structures in the farms. As leakage current is often voltage-dependent, one must take care to apply the best approximation of voltages the different FETs will experience during device measurement when performing these leakage measurements for subtraction. The sense lines are designed for minimal leakage such that leakage due to the disabled pull-down/-up FETs are able to compensated. Pass FET gate leakages may vary slightly when performing actual device measurements compared to open structure measurements due to voltage differential across the pass gate when carrying current, but discrepancies are simulated to be a very small component of the total leakage and do not impede the extraction of model parameters.
@@ IV Results
Figure 3: (a) Drain current of a single NFET as a function of gate-source bias ($I_{d}$-$V_{gs}$ curves). Measurements were taken at $V_{ds}$ of 0.05 V with backgate voltages ($V_{bs}$) of 0 V, 1 V and 2 V (increasing backgate voltage corresponding to lighter shade). Measurements taken at 300 K and 4 K can be seen with solid and dashed lines, respectively. (b) $I_{d}$-$V_{gs}$ curves at $V_{ds}=0.05$ V and backgate voltage of 0 V at 4 K for four geometries of NFETs represented by different colors. Dashed lines represent measurement data, while solid lines represent BSIM-IMG v102.9.6 model simulations using model parameters extracted from the 4 K measurement data demonstrating a relatively good fit as a proof-of-concept for cryo-PDK model building. (c) $I_{d}$-$V_{gs}$ curve (solid line) at $V_{ds}=0.05$ V and backgate voltage of 2 V at 4 K for an NFET demonstrating the intersubband scattering effect. The corresponding transconductance ($g_{m}$) curve is smoothed and plotted as a dashed line to further display the characteristic behavior of this effect.
@@ IV-A Representative I-V Curves
Figure 4: (a) Depiction of a die containing device farms and integrated multiplexer, highlighting the location of one of the farms. Each farm consists of a 2D array of transistors of various geometries in a pseudo-random common-centroid configuration. The relative position of PFET devices with 2 nominal geometries are represented by red and blue squares (smaller and larger dimensions, respectively). (b) Parameter spread of threshold voltage ($V_{th}$) and on current $I_{dON}$ ($V_{gs}=-0.8$ V and $V_{ds}=-0.8$ V) of devices with the corresponding color in (a). The geometry represented by the blue points has a greater width-to-length ratio, explaining the greater magnitude of $I_{dON}$. Measurements were obtained across two nominally identical dies (indicated by square and asterisk point markers. Insets (c) and (d) each highlight the parameter spread of devices within a single nominal geometry.
@@ IV-B Parameter Selection and Sweep Rates
For DC I-V characterization at cryogenic temperatures, the optimal step size within a voltage sweep considers the trade-off between measurement time and the accuracy in an extracted model’s I-V curve shape. As a rough estimate, $N_{gs}\approx 40-100$ points are typically required for a $V_{gs}$ sweep extending from the subthreshold region into the strong inversion regime to provide sufficient information for model building; fewer points are generally required for $V_{ds}$ sweeps. Careful determination of threshold voltage and transition between different regions of operation is crucial for model parameter extraction and validation. In these regions in which the real device effects of channel length modulation, drain-induced threshold shifts, etc. [20] can be observed, smaller steps may be justified. In the deep-subthreshold region, FET devices also show gate-induced drain- and source-lowering effects [21]. Due to the very low current values, modelling such regions of operation requires denser sweep points. However, the kind of circuit to be designed (e.g. digital, high-power vs. low-power analog, RF) must be considered. For some applications, very low current regions of operation may not require a very accurate model and therefore a coarse sweep for this region of operation will suffice. For digital circuits the threshold voltage and the current in saturation are critical, so step size may be decreased in those regions of operation to ensure an accurate model.
@@ IV-C Bulk Measurements
To support high-quality model-building, $V_{gs}$ sweeps must be conducted at several $V_{ds}$ bias voltages, and $V_{ds}$ sweeps are often conducted at a few $V_{gs}$ bias voltages. Then this entire set of $V_{gs}$ and $V_{ds}$ sweeps are carried out at multiple backgate bias voltages. Thus, we can estimate total measurement time for a farm as a multiplication of several key chosen parameters as given in Table I. The total number of measurement points per device is of order $10^{3}$. For a farm of 1024 devices, this gives a total of $N_{pts}\approx 10^{6}$ data points from a farm. Using the values in Table I gives a minimum raw measurement time of at least 7.5 hours not including any overheads. The total time taken for collection of a full measurement set is greatly increased by a characterized settling time requirement of $\geq 20$ ms. Accounting for this set-measure delay along with instrument communication time related to multiplexer selection, actual measurement time is closer to 20 hours. Optimising the multiplexer selection time (e.g. via digital pattern pre-loading) would make only a small difference, reducing the total time to a minimum of $\sim 19$ hours. A better route to reducing total measurement time is to reduce the required set-measure delay, which requires further optimisation of wiring both on-chip and in the probe. Additionally, if a sufficient number of SMUs are available, a device in each of the four farms on a die may be measured simultaneously in parallel, drastically reducing the effective measurement time per device. However, one must consider the effects cross-device heating on the same die may be significant. Furthermore, multiple dies may be connected in such a way that they share power supply lines and digital control signals (but separate analog connections to their farms) to allow parallel measurements across several dies at once, providing further potential orders of efficiency improvement.
@@
TABLE I: Representative time budget for a full measurement set across a 1024 device transistor farm.
Chosen Parameters Quantity Unit
@@
Integration time 1 PLC
Total raw measurement time 7.5 h
Typical delay sources delay units
@@ VI Acknowledgements
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may be no longer accessible.

### v2

# Method for efficient large-scale cryogenic characterization of CMOS technologies

@@ Abstract
Index Terms:
 Cryogenic Electronics, Cryo-CMOS, Semiconductor Device Modeling, I-V Curves, MOSFET, Silicon-on-Insulator, Quantum Computing.
@@ I Introduction
Quantum computing devices based in semiconductor systems offer many possible advantages of ready scalability by harnessing existing fabrication technology [1]. Application-specific integrated circuits (ASICs) fabricated using established foundry processes can be operated at low temperature to support the scalable control of the large number of qubits necessary for error correction [2]. Integrated multiplexing and signal processing reduces the number of connecting lines between cryogenic stages and room-temperature instruments and reduces the complexity of the room-temperature readout instrumentation required [3]. Significant progress has been made in the development of quantum-orientated CMOS-based cryogenic control systems [4, 5, 6, 7] which can be utilised in a variety of quantum device platforms including superconducting qubits [8, 9] and trapped ions [10].
At the most fundamental level, confident cryogenic integrated circuit design begins with test and validation of the elementary transistor components in the technology node under realistic cryogenic operation conditions. This is necessary because the operating parameters of silicon devices are dependent on temperature, and this dependence cannot be extrapolated from foundry-provided models (often only rated down to -55 °C [11]) due to the onset of cryogenic effects which contribute significantly to very low-temperature behavior. The device parameters captured in the process design kit (PDK) used in electronic design automation (EDA) tools need to accurately reflect realistic operation at low temperature. Significant efforts to characterize and model the behavior of low-temperature transistors has been performed [12, 13, 14, 15].
@@
Capturing a comprehensive set of cryogenic test data presents logistical challenges. The overhead of wire-bonding individual transistor device terminals combined with practical constraints of die size, cryostat wiring, and thermal cycle times make this approach impractical. At room temperature, these measurements are typically performed at the wafer scale, using an automated wafer prober to access dedicated pads for each transistor across large arrays of devices. There are advantages to this approach. Wafer-level probing of individual devices allows for simpler and more standardized de-embedding approaches to be followed. It has been shown that this method can be extended to cryogenic environments by utilizing highly specialized cryogenic probe systems [17]. This method allows wafer-scale variability to be measured in an analogous way to the traditional foundry PDK development process; however, it lacks flexibility and accessibility due to the heavy dependence on expensive, specialist equipment and inefficient use of wafer area.
@@ III Measurement System
Figure 1: (a) Block diagram of the key components in the measurement system. A die containing multiple device arrays, each consisting of 1024 transistors accessed via a multiplexer is wire-bonded to a demount-able PCB "daughter card." A corresponding multi-pin socket mounted to a liquid helium dipping probe provides an interface to the digital circuitry necessary to configure the state of the multiplexer as well as the analog circuity used to conduct the measurements on the selected device. (b) Photograph of the bonded die. By recessing the die within a multi-level PCB and utilizing the surface of both levels, the density of bond pads can be increased, minimizing the length and approach angle of bond wires needed.
@@ III-B Integrated Circuits (DUTs) and Sample Packaging
The core multiplexer unit cell structure contains an analog signal pass gate, or switch, formed by an N-channel FET (NFET) and P-channel FET (PFET) in parallel with digitally driven gate voltages to define the on and off states as shown in Figure 2(a). Since the increased threshold voltages of these devices at cryogenic temperatures result in substantially increased on-resistance when the pass voltage is near the midpoint of the high and low supply voltages, $V_{DD}$ and $V_{SS}$ (in some cases both the NFET and PFET may be in the subthreshold regime even with the pass gate in the "on" state), the backgate voltages (not shown) of the multiplexer pass gate FETs are externally driven to enable shifting the device thresholds nearer to their room-temperature values [22].
@@ III-E Analog Instrumentation
Figure 2: (a) High-level circuit diagram of an on-die multiplexer structure allowing analogue connections to be established with Kelvin sense to each terminal of a selected transistor. Gate contacts are established at opposite sides of the channel to allow gate resistance thermometry to be performed. A pair of 32-bit registers (one for one-hot row selection and another for one-hot column selection) are used to turn on the pass gate to the device at the selected row/column pair. Each signal pass gate is formed by an NFET and PFET in parallel. For a farm containing NFET devices for measurement as shown, a pull-down FET is activated for devices that are not selected; for a farm containing PFET devices for measurement, a pull-up FET is used instead. (b) Diagram of analog instrument connections used to conduct a 6-terminal measurement with Kelvin sense (FH, SH, FL, and SL corresponding to force-high, sense-high, force-low, and sense-low channels of SMUs, respectively). The ESD protection diodes in blue demonstrate the value of using SMU3 as a reference to the system ground. SMU3 in this case can bias the source at an arbitrary reference potential with respect to system ground, enabling full control and optimization of measurements to avoid excess forward-bias currents flowing in any ESD diodes while maintaining high-impedance sense lines.
Analog transistor DC characteristics can be probed with multiple source-measure units (SMUs) as shown in Figure 2(b). Typically, one unit controls drain-source bias and the other controls gate-source bias. These are bound to a common reference point whose potential can be arbitrarily set with respect to chassis ground by the third SMU. When using a Kelvin sensing configuration, each device terminal must have independent lines for current injection ("force") and voltage monitoring ("sense"). These can be implemented as parallel switching routes to a multiplexed device terminal as previously described. Sensing voltages on the device side of the multiplexers allows for accurate real-time compensation for voltage drops due to the impedance of the measurement leads and multiplexer switches. This system also gives control over how the analog circuitry as a whole is referenced to the low-side supply $V_{SS}$. This allows the full range of accessible potentials on each terminal, defined by ESD protection structures, to be utilized.
@@ III-F Leakage Subtraction
Figure 3: Example 0.8-V farm multiplexer switch characteristics measured at 4 K across a range of pass voltages (relative to the ground voltage of the switch circuitry) with appropriate switch FET backgate biases applied to enable cryogenic operation. (a) Leakage current measurement conducted on an open structure, including leakage current effects from both the force line switch and the connected sense line switch. (b) Force line and sense line switch on-resistances independently measured at 1 µA pass current.
@@
Leakage on the die arises mainly from ESD protection structures, off-state drain-source leakage through the disabled pass FETs of other multiplexer channels, off-state leakage through the disabled pull-down/-up FETs, and gate leakage from the various FETs. These leakage effects can be measured (and then subtracted from actual device measurements) by measuring the open structures in the farms. As leakage current is often voltage-dependent, one must take care to apply the best approximation of voltages the different FETs will experience during device measurement when performing these leakage measurements for subtraction. The sense lines and their switches are designed for minimal leakage such that leakage due to the disabled pull-down/-up FETs are able to compensated. An example open structure leakage current measurement (including leakage from both force and sense switches) along with separate force and sense switch on-resistances are plotted across switch pass voltage in Figure 3.
@@ IV Results
Figure 4: (a) Drain current of a single NFET as a function of gate-source bias ($I_{d}$-$V_{gs}$ curves). Measurements were taken at $V_{ds}$ of 0.05 V with backgate voltages ($V_{bs}$) of 0 V, 1 V and 2 V (increasing backgate voltage corresponding to lighter shade). Measurements taken at 300 K and 4 K can be seen with solid and dashed lines, respectively. (b) $I_{d}$-$V_{gs}$ curves at $V_{ds}=0.05$ V and backgate voltage of 0 V at 4 K for four geometries of NFETs represented by different colors. Dashed lines represent measurement data, while solid lines represent BSIM-IMG v102.9.6 model simulations using model parameters extracted from the 4 K measurement data demonstrating a relatively good fit as a proof-of-concept for cryogenic PDK model building. (c) $I_{d}$-$V_{gs}$ curve (solid line) at $V_{ds}=0.05$ V and backgate voltage of 2 V at 4 K for an NFET demonstrating the intersubband scattering effect. The corresponding transconductance ($g_{m}$) curve is smoothed and plotted as a dashed line to further display the characteristic behavior of this effect.
@@ IV-B Parameter Selection and Sweep Rates
For DC I-V characterization at cryogenic temperatures, the optimal step size within a voltage sweep considers the trade-off between measurement time and the accuracy in an extracted model’s I-V curve shape. As a rough estimate, measurements at approximately $40-100$ gate-source biases are typically required for a $V_{gs}$ sweep extending from the subthreshold region into the strong inversion regime to provide sufficient information for model building; fewer points are generally required for $V_{ds}$ sweeps. Careful determination of threshold voltage and transition between different regions of operation is crucial for model parameter extraction and validation. In these regions in which the real device effects of channel length modulation, drain-induced threshold shifts, etc. [31] can be observed, smaller steps may be justified. In the deep-subthreshold region, FET devices also show gate-induced drain- and source-lowering effects [32]. Due to the very low current values, modelling such regions of operation requires denser sweep points. However, the kind of circuit to be designed (e.g. digital, high-power vs. low-power analog, RF) must be considered. For some applications, very low current regions of operation may not require a very accurate model and therefore a coarse sweep for this region of operation will suffice. For digital circuits the threshold voltage and the current in saturation are critical, so step size may be decreased in those regions of operation to ensure an accurate model.
@@ IV-C Bulk Measurements
Figure 5: (a) Depiction of a die containing device farms and integrated multiplexer, highlighting the location of one of the farms. Each farm consists of a 2D array of transistors of various geometries in a pseudo-random common-centroid configuration. The relative position of PFET devices with 2 nominal geometries are represented by red and blue squares (smaller and larger dimensions, respectively). (b) Parameter spread at $V_{bs}=-0.8$ V of threshold voltage ($V_{th}$) and on current $I_{dON}$ ($V_{gs}=-0.8$ V and $V_{ds}=-0.8$ V) of devices with the corresponding color in (a). The geometry represented by the blue points has a greater width-to-length ratio, explaining the greater magnitude of $I_{dON}$. Measurements were obtained across two nominally identical dies (indicated by square and asterisk point markers. Insets (c) and (d) each highlight the parameter spread of devices within a single nominal geometry.
@@
Statistical data for each distribution in Figure 5 can be found along with comparisons to their simulated room-temperature counterparts in Table I. While the level of variability observed in the smaller geometry is much greater than that of the large geometry, it is actually the larger geometry whose relative increase in variability compared to room-temperature values is more significant. In fact, for one of the dies, the observed cryogenic variability of the smaller geometry (which represents very nearly the minimum drawn dimensions available for a PFET in the studied technology node) is actually lower than its predicted room-temperature variability. This underscores the need for large sample sizes of devices and multiple dies to be characterized to draw accurate conclusions about population variability. The relative (percent) increases in parameter sample standard deviations from room temperature to 4 K are overall comparable to published cryogenic variability increases in other technology nodes [14, 16]. This increase in variability is critical to characterize as part of cryogenic PDK development to allow circuit designers to make informed decisions about performance and variability trade-offs and to determine what kind of trim capabilities may be needed to meet desired performance specifications for a given circuit.
TABLE I: Sample statistics of extracted parameters for selected PFET geometries at 4 K with room-temperature comparisons
Larger Larger Smaller Smaller
geometry, geometry, geometry, geometry,
Parameter die 1 die 2 die 1 die 2
Sample size 32 32 32 32
4K $\overline{V_{th}}$ (mV) -378.2 -393.7 -420.2 -410.0
4K $\sigma(V_{th})$ (mV) 4.3 4.2 27.1 33.1
$\underset{\textrm{RT}\rightarrow\textrm{4K}}{\Delta}[\overline{V_{th}}]$ (mV)* -135.6 -151.0 -141.8 -131.6
$\underset{\textrm{RT}\rightarrow\textrm{4K}}{\Delta}[\sigma(V_{th})]$ (%)* 66.5% 63.8% 2.7% 25.3%
4K $\overline{I_{dON}}$ (µA) -403.1 -387.5 -32.2 -34.1
4K $\sigma(I_{dON}/\overline{I_{dON}})$ (%) 1.7% 1.4% 7.9% 13.4%
$\underset{\textrm{RT}\rightarrow\textrm{4K}}{\Delta}[\overline{I_{dON}}]$ (%)* -1.0% -4.9% -18.3% -13.5%
$\underset{\textrm{RT}\rightarrow\textrm{4K}}{\Delta}[\sigma(I_{dON}/\overline{I_{dON}})]$ (%)* 74.8% 43.1% -17.6% 39.8%
*Delta parameters represent 4-K measurements’ difference from room-temperature Monte Carlo simulations of local mismatch variation (sample size of 100) using foundry-provided model cards and taking into account local layout effects and parasitic extraction for the devices in the farm.
@@
To support high-quality model-building, $V_{gs}$ sweeps must be conducted at several $V_{ds}$ bias voltages, and $V_{ds}$ sweeps are often conducted at a few $V_{gs}$ bias voltages. Then this entire set of $V_{gs}$ and $V_{ds}$ sweeps are carried out at multiple backgate bias voltages. Thus, we can estimate total measurement time for a farm as a multiplication of several key chosen parameters as given in Table II. The total number of measurement points per device is of order $10^{3}$. For a farm of 1024 devices, this gives a total of approximately $10^{6}$ data points per farm. Using the values in Table II gives a minimum raw measurement time of at least 7.5 hours not including any overheads. The total time taken for collection of a full measurement set is greatly increased by a characterized settling time requirement of $\geq 20$ ms. Accounting for this set-measure delay along with instrument communication time related to multiplexer selection, actual measurement time is closer to 20 hours. Optimising the multiplexer selection time (e.g. via digital pattern pre-loading) would make only a small difference, reducing the total time to a minimum of $\sim 19$ hours. A better route to reducing total measurement time is to reduce the required set-measure delay, which requires further optimisation of wiring both on-chip and in the probe. Additionally, if a sufficient number of SMUs are available, a device in each of the four farms on a die may be measured simultaneously in parallel, drastically reducing the effective measurement time per device. However, one must consider the effects of cross-device heating on the same die may be significant. Furthermore, multiple dies may be connected in such a way that they share power supply lines and digital control signals (but separate analog connections to their farms) to allow parallel measurements across several dies at once, providing further potential orders of efficiency improvement.
@@
TABLE II: Representative time budget for a full measurement set across a 1024 device transistor farm.
Parameter Quantity Unit
@@
Integration time 20 ms